OpenStudy (anonymous):

A latter is propped up 3 feet away from the side of a building. The length of the latter is 8ft. What is the height of the building?

OpenStudy (kira_yamato):



OpenStudy (kira_yamato):

Using the Pythagorean Theorem, \[h^2 = 8^2 - 3^2\]\[h = \sqrt{55}\]